OpenStudy (anonymous):

choose the ordered pair (x,y) that is a solution of the equation. -18x+9y=54 answers; (7,14) (3,12) (2,9) (14,13)

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

-18x + 9y = 54 time to sub again... (7,14) -18(7) + 9(14) = 54 -126 + 126 = 54 0 = 54 (not this one) (3,12) -18(3) + 9(12) = 54 -54 + 108 = 54 54 = 54 (correct) Its this one :)